OpenStudy (anonymous):

Which one is a rule for the nth term of the sequence 3, 15, 75, 375...?

OpenStudy (anonymous):

A. an = 3(5)n − 1B. an = 5(3)n − 1C. an = 3(5)n + 1D. an = 5(3)n + 1

OpenStudy (mertsj):

\[a _{n}=a _{1}r ^{n-1}\]